### v2

# Post-Minkowskian Theory Meets the Spinning Effective-One-Body Approach for Bound-Orbit Waveforms

Abstract
Driven by advances in scattering amplitudes and worldline-based methods, recent years have seen significant progress in our ability to calculate gravitational two-body scattering observables. These observables effectively encapsulate the gravitational two-body problem in the weak-field and high-velocity regime (post-Minkowskian, PM), with applications to the bound two-body problem and gravitational-wave modeling. We leverage PM data to construct a complete inspiral-merger-ringdown waveform model for non-precessing spinning black holes within the effective-one-body (EOB) formalism: SEOBNR-PM. This model is closely based on the highly successful SEOBNRv5 model, used by the LIGO-Virgo-KAGRA Collaboration, with its key new feature being an EOB Hamiltonian derived by matching the two-body scattering angle in a perturbative PM expansion. The model performs remarkably well, showing a median mismatch against 441 numerical-relativity (NR) simulations that is somewhat lower than a similarly calibrated version of SEOBNRv5. Comparisons of the binding energy with NR also demonstrate better agreement than SEOBNRv5, despite the latter containing additional calibration to NR simulations.
preprint: HU-EP-24/16-RTG
Introduction – Since the initial detection of a gravitational wave (GW) from a binary-black-hole (BBH) merger Abbott et al. (2016a), the LIGO-Virgo-KAGRA Collaboration Aasi et al. (2015); Acernese et al. (2015); Akutsu et al. (2021) and independent analyses have identified about 100 mergers of compact binaries Abbott et al. (2017a, 2023a); Nitz et al. (2023); Olsen et al. (2022); Mehta et al. (2023); Wadekar et al. (2023). These observations have begun to reveal the distributions of BH masses and spins Abbott et al. (2023b), improved constraints on the neutron-star equation of state Abbott et al. (2018a), obtained independent measurements of the Hubble-Lemaître parameter Abbott et al. (2017b, 2023c), and validated General Relativity Abbott et al. (2016b, 2021a, 2021b).
Enhancements in the sensitivity of current GW detectors, coupled with the development of next-generation observatories like the Einstein Telescope and Cosmic Explorer Punturo et al. (2010); Kalogera et al. (2021); Reitze et al. (2019), as well as future space-based detectors such as LISA Amaro-Seoane et al. (2017), TianQin Luo et al. (2016) or Taiji Wu et al. (2021), are poised to dramatically increase the number of detectable GW sources. These advancements will enable observations with a signal-to-noise ratio up to two orders of magnitude higher than what is currently achievable Borhanian and Sathyaprakash (2024), necessitating a commensurate improvement in the accuracy of waveform models. Recent research Dhani et al. (2024) has demonstrated that even state-of-the-art waveform models, designed for quasi-circular, spin-precessing BBHs, exhibit systematic biases when applied to future LIGO-Virgo-KAGRA runs and next-generation detectors. This bias becomes pronounced, especially for high-spin rates and significant asymmetries in spins and masses. Addressing the challenge of waveform accuracy is essential to realizing the full scientific potential of future runs and detectors Abbott et al. (2018b); Evans et al. (2021); Bogdanov et al. (2022); Borhanian and Sathyaprakash (2024); Colpi et al. (2024) and avoiding false claims of General Relativity violations Toubiana et al. (2024); Gupta et al. (2024).
Waveform models for compact binaries are crafted by synergistically combining analytical and numerical relativity (NR) results. NR tackles the formidable task of solving Einstein’s equations on supercomputers Pretorius (2005); Campanelli et al. (2006); Baker et al. (2006), a process notorious for its time-intensive nature. On the other hand, perturbative methods are used to obtain approximate solutions to Einstein’s equations, offering analytic formulas that are swift to evaluate. Three primary perturbative approaches have been developed: post-Newtonian (PN) theory Futamase and Itoh (2007); Blanchet (2014); Porto (2016); Schäfer, Gerhard and Jaranowski, Piotr (2018); Levi (2020); Jaranowski and Schaefer (1998); Damour, Thibault and Jaranowski, Piotr and Schäfer, Gerhard (2014); Jaranowski and Schäfer (2015); Bernard et al. (2016, 2017); Damour, Thibault and Jaranowski, Piotr and Schäfer, Gerhard (2016); Blanchet et al. (2023a, b); Foffa and Sturani (2019); Foffa et al. (2019); Blümlein, J. and Maier, A. and Marquard, P. and Schäfer, G. (2020) applicable in the weak-field and small-velocity limit, post-Minkowskian (PM) theory Westpfahl and Goller (1979); Westpfahl and Hoyler (1980); Bel et al. (1981); Westpfahl (1985); Schäfer (1986); Ledvinka et al. (2008); Buonanno et al. (2022); Travaglini et al. (2022); Bjerrum-Bohr et al. (2022); Kosower et al. (2022) in the weak-field regime, and the gravitational-self force (GSF) formalism Mino et al. (1997); Quinn and Wald (1997); Barack et al. (2002); Barack and Ori (2003); Gralla and Wald (2008); Detweiler (2008); Keidl et al. (2010); van de Meent (2018); Pound (2012); Pound et al. (2020); Gralla and Lobo (2022); Pound and Wardell (2021); Warburton et al. (2021); Wardell et al. (2023) for the small mass-ratio limit. These analytical results are then synthesized in the effective-one-body (EOB) approach Buonanno and Damour (1999, 2000); Damour et al. (2000); Damour (2001); Buonanno et al. (2006), which efficiently resums the perturbative calculations for the inspiral while retaining known non-perturbative results for BHs, achieving high accuracy for current observing runs via calibration to NR Buonanno et al. (2007a, b); Damour and Nagar (2008); Pan et al. (2011a); Damour et al. (2013); Pan et al. (2014); Taracchini et al. (2014); Bohé et al. (2017); Babak et al. (2017); Nagar et al. (2018); Ossokine et al. (2020); Gamba et al. (2022); Nagar et al. (2023); Pompili et al. (2023); Ramos-Buades et al. (2023).
Thus far, the EOB waveform models utilized by the LIGO-Virgo-KAGRA Collaboration [^1] The LIGO-Virgo-KAGRA Collaboration also utilizes phenomenological frequency- and time-domain waveform models, which are built by combining PN, EOB, and NR results Ajith et al. (2007); Pratten et al. (2021); Estellés et al. (2022). Additionally, it employs waveforms directly interpolated from NR simulations Blackman et al. (2017a); Varma et al. (2019a), where available. have primarily relied on resummations of the PN expansion, with the exception of Refs. Ramos-Buades et al. (2023); van de Meent et al. (2023), which included second-order GSF results Warburton et al. (2021) for the gravitational modes and radiation-reaction force. Given recent advancements in PM Damour (2016); Cheung et al. (2018); Bern et al. (2019, 2023, 2024a); Kosower et al. (2019); Bjerrum-Bohr et al. (2018); Cristofoli et al. (2019); Damgaard et al. (2019); Brandhuber et al. (2021); Vines (2018); Kälin, Gregor and Porto, Rafael A. (2020); Kälin, Gregor and Liu, Zhengwen and Porto, Rafael A. (2020); Mogull et al. (2021); Jakobsen et al. (2022a, b); Bhattacharyya et al. (2024) and GSF Warburton et al. (2021); Wardell et al. (2023), there is now significant interest in exploring and developing waveform models that combine information from various perturbative methods in innovative ways. The aim is to address the waveform-accuracy challenge. In this regard, the PM approach is particularly interesting since an (n + 1) PM-order ($G^{n+1}$) Hamiltonian includes all information up to the nPN order (wherein $v^{2}/c^{2}\sim GM/(rc^{2})$), and additional weak-field, high-velocity information from infinitely higher PN orders, making it suitable for systems with high velocities or large eccentricities at fixed periastron distances Khalil et al. (2022). Using sophisticated quantum-field-theory-based methods, tremendous progress has been made on the precision PM frontier using both scattering amplitudes Bern et al. (2021a, 2022a); Damgaard et al. (2023) and worldlines Dlapa et al. (2022, 2023a); Jakobsen et al. (2023a, 2024). This progress is largely due to a blend of a clever and efficient organization of perturbative calculations and formal mathematical developments in understanding the properties of multi-loop integrals Parra-Martinez et al. (2020); Weinzierl (2022); Abreu et al. (2022); Blümlein and Schneider (2022); Frellesvig et al. (2024a, b); Klemm et al. (2024). These developments were primarily driven over the past several decades to address precision-collider physics. Furthermore, several sophisticated techniques, such as generalized unitarity Bern et al. (1995); Bern and Huang (2011), double copy Bern et al. (2010, 2024b, 2022b); Adamo et al. (2022), supersymmetry Jakobsen et al. (2022a) and massive higher spins Chung et al. (2019); Arkani-Hamed et al. (2020); Bern et al. (2021b); Bautista et al. (2023a, b); Chiodaroli et al. (2022); Aoude et al. (2023); Cangemi et al. (2024) have also been used to further enhance these computations.
In this Letter, leveraging on Refs. Khalil et al. (2023); Pompili et al. (2023); Buonanno et al. (2024), we present the first PM-informed spinning EOB waveform model: SEOBNR-PM, encompassing the inspiral, as well as the merger and ringdown phases. This model incorporates the most recent findings from PM theory into the EOB Hamiltonian, and it is mildly calibrated to NR waveforms. The SEOB-PM Hamiltonian (so named as it does not include an NR calibration term) includes the non-spinning (conservative) 4PM Bern et al. (2022a); Dlapa et al. (2022) and spinning 5PM terms Guevara et al. (2019); Bern et al. (2021b); Kosmopoulos and Luna (2021); Chen et al. (2022); Jakobsen and Mogull (2022, 2023); Febres Cordero et al. (2023); Jakobsen et al. (2023a), alongside the known non-spinning 4PN Damour, Thibault and Jaranowski, Piotr and Schäfer, Gerhard (2014); Jaranowski and Schäfer (2015); Bernard et al. (2016, 2017); Damour, Thibault and Jaranowski, Piotr and Schäfer, Gerhard (2016) contributions, which also corrects the tails from unbound to bound orbits up to that order. [^2] We plan to update our SEOBNR-PM model in the near future to include the recent findings in Ref. Dlapa et al. (2024), which enables derivation of the local-in-time 4PM EOB Hamiltonian for bound orbits. Our PM counting is a physical one, with spin orders contributing in addition to loop orders (see Table II in Ref. Buonanno et al. (2024)). We construct our SEOBNR-PM model within the pySEOBNR code Mihaylov et al. (2023), which was recently built to make the development of SEOBNR models, including the calibration to NR waveforms, highly efficient. As an example, we show in Fig. 1 the agreement between the (mildly) calibrated SEOBNR-PM and NR for a spinning BBH coalescence.
Figure 1: An SEOBNR-PM inspiral-merger-ringdown waveform generated using the pySEOBNR code Mihaylov et al. (2023), compared against the NR simulation SXS:BBH:1445, after a low-frequency alignment by a time and phase shift. The time $t=0$ corresponds to the peak of the $(2,2)$ mode of the NR waveform.
EOB framework for waveforms – We use geometric units $G=1=c$, and set $M=m_{1}+m_{2}$ and $\nu=m_{1}m_{2}/M^{2}$, where $m_{1}$ and $m_{2}$ are the BH’s masses. In the EOB formalism, the binary’s conservative dynamics is described by the EOB Hamiltonian $H_{\rm EOB}=M\,\sqrt{1+2\nu(H_{\rm eff}/\mu-1)}$, where $H_{\rm eff}$ is the Hamiltonian of an effective test body of mass $\mu=\nu\,M$ moving in the (deformed) Kerr spacetime, with $40\leq\nu\leq 1/$ 4 the deformation parameter. We also introduce the mass ratio $q=m_{1}/m_{2}>1$. We limit to nonprecessing spins (i.e., aligned spins) and introduce the spin lengths $a_{i}=m_{i}\chi_{i}$, with $a_{\pm}=M\chi_{\pm}=a_{1}\pm a_{2}$.
In the center-of-mass frame, the inspiral-plunge dynamics, for aligned-spin BHs, is computed from the EOB equations of motion Buonanno and Damour (2000); Pan et al. (2011a):
$\dot{r}$ $=\frac{\partial H_{\rm EOB}}{\partial p_{r}}\,,$ $\dot{p}_{r}$ $=-\frac{\partial H_{\rm EOB}}{\partial r}+\frac{p_{r}}{p_{\phi}}\mathcal{F}_{\phi}\,,$ (1a)
$\dot{\phi}$ $=\frac{\partial H_{\rm EOB}}{\partial p_{\phi}}\,,$ $\dot{p}_{\phi}$ $=\mathcal{F}_{\phi}\,,$ (1b)
where ($r$,$\phi$,$p_{r}$,$p_{\phi}$) are the canonical variables in polar coordinates. (The construction of $H_{\rm eff}$ will be described in the next section.) Employing results from the SEOBNRv5 model Khalil et al. (2023); Pompili et al. (2023), the radiation-reaction force ($\mathcal{F}_{\phi}$) is computed by summing over the PN GW modes (augmented with GSF information van de Meent et al. (2023)) in a factorized form Damour et al. (2009); Damour and Nagar (2007); Pan et al. (2011b); Damour and Nagar (2008); Pompili et al. (2023), which are used to obtain the inspiral-plunge modes after enhancing them by non-quasi-circular corrections Damour et al. (2003); Taracchini et al. (2014); Bohé et al. (2017); Pompili et al. (2023) during the plunge.
For the merger-ringdown part of the EOB waveform, we use instead a phenomenological ansatz Damour and Nagar (2014); Bohé et al. (2017); Cotesta et al. (2018); Pompili et al. (2023), informed by NR and BH perturbation theory, as realized in the SEOBNRv5 model Pompili et al. (2023). The start of the merger-ringdown waveform is enforced to be at the peak of the $(2,2)$-mode amplitude. The gravitational polarizations can be written as $h_{+}-ih_{\times}=\sum_{\ell,m}{}_{-2}Y_{\ell m}(\varphi,\iota)h_{\ell m}(t)$, where ${}_{-2}Y_{\ell m}(\varphi,\iota)$ are the -2 spin-weighted spherical harmonics, with $\varphi$ and $\iota$ being the azimuthal and polar angles to the observer, in the source frame. In the EOB approach, the inspiral-merger-ringdown $(\ell,m)$ modes are given by
$h_{\ell m}=\begin{cases}h_{\ell m}^{\rm insp-plunge},&t<t_{\text{peak }}^{22},\\ h_{\ell m}^{\rm merg-RD},&t>t_{\text{peak }}^{22}.\end{cases}$ (2)
where $t_{\text{peak }}^{22}$ is the time at which the (2,2) mode has a peak, generally associated to the merger time. Such a time is suitably chosen to agree with the corresponding time in NR waveforms (see below).
PM-informed EOB Hamiltonian – We employ an effective Hamiltonian similar to that recently introduced in the SEOB-PM scattering model Buonanno et al. (2024):
$H_{\rm eff}=\frac{Mp_{\phi}(g_{a_{+}}a_{+}+g_{a_{-}}\delta a_{-})}{r^{3}+a_{+}^{2}(r+2M)}$ (3)
$\qquad+\sqrt{A\left(\mu^{2}+\frac{p_{\phi}^{2}}{r^{2}}+(1+B_{\rm np}^{\rm Kerr})p_{r}^{2}+B_{\rm npa}^{\rm Kerr}\frac{p_{\phi}^{2}a_{+}^{2}}{r^{2}}\right)}\,,$
where $\delta=(m_{1}-m_{2})/M$, while $B_{\rm np}^{\rm Kerr}=\chi_{+}^{2}u^{2}-2u$ and $B_{\rm npa}^{\rm Kerr}=-{(1+2u)}/{[r^{2}+a_{+}^{2}(1+2u)]}$, where $u=M/r$ is the dimensionless PM counting parameter. In the probe limit $\nu\to 0$, $H_{\rm eff}$ reduces to the Hamiltonian of a probe $\mu$ moving under the influence of a Kerr BH with mass $M$ and directed spin length $a_{+}$. This Hamiltonian is determined by computing the scattering angle and matching it to established PM results, but here we use only the conservative part of the angle containing terms with even powers in the center-of-mass momentum $p_{\infty}=\mu\sqrt{\gamma^{2}-1}$, where $\gamma=E_{\rm eff}/\mu>1$ for scattering trajectories.
Following Ref. Buonanno et al. (2024), the $\nu$-corrections with respect to the probe limit are built into the $A$-potential and the gyro-gravitomagnetic factors as $A={(1-2u+\chi_{+}^{2}u^{2}+\Delta A)}/{[1+\chi_{+}^{2}u^{2}(2u+1)]}$ and $g_{a_{\pm}}={\Delta g_{a_{\pm}}/}{u^{2}}$. These respectively carry the even- and odd-in-spin corrections, and are PM expanded up to a physical 5PM ($u^{5}$) order (see Table II in Ref. Buonanno et al. (2024)):
$\Delta A$ $=\sum_{n=2}^{5}u^{n}\Delta A^{(n)}+\Delta A^{\rm 4PN}\,,$ $\Delta g_{a_{\pm}}$ $=\sum_{n=2}^{5}u^{n}\Delta g_{a_{\pm}}^{(n)}\,.$ (4)
The $\gamma$-dependent coefficients $\Delta A^{(n)}$ and $\Delta g_{a_{\pm}}^{(n)}$ are series expanded in even powers of the spins, up to a highest quartic order at 5PM. We lack an analytic 5PM term only in the non-spinning case, where the complete result is not currently known (see Ref. Driesse et al. (2024) for the recently derived 1GSF conservative contribution). Technically, as $\gamma=E_{\rm eff}/\mu\equiv H_{\rm eff}/\mu$, the Hamiltonian (3) is self-dependent. To produce an expression depending only on the canonical variables ($r$,$p_{r}$,$p_{\phi}$), we interpret $\gamma=H_{\rm Kerr}/\mu$ within these deformations, plus whatever corrections are required in order to ensure the full Hamiltonian $H_{\rm EOB}$ is correct up to the desired PM order. This procedure was used previously in the non-spinning case Damour (2018); Antonelli et al. (2019); Khalil et al. (2022), and is fully described in the Supplemental Material.
An important subtlety within our Hamiltonian is the presence of non-local-in-time contributions (tails). These imply a dependence on the full past history of the binary, and thus distinguish between elliptic and hyperbolic (scattering) trajectories. In the scattering Hamiltonian presented in Ref. Buonanno et al. (2024), tails are signaled by factors of $\log(\gamma^{2}-1)$, which develops an imaginary part when $\gamma<1$ for bound orbits. To produce a real Hamiltonian, we therefore replace $\log(\gamma^{2}-1)\to\log(u)$ (see Supplemental Material for details). We also include the 4PN non-spinning bound-orbit correction $\Delta A^{\rm 4PN}$ in Eq. (4),
$\Delta A^{\rm 4PN}$ $=u^{4}(\gamma^{2}-1)c_{1}+u^{5}(c_{2}+c_{3}\log u)\,,$ (5)
ensuring the correct bound-orbit dynamics at 4PN order in the non-spinning case (the numerical coefficients $c_{i}$ are provided in the Supplemental Material). We verify our complete EOB Hamiltonian up to 4.5PN order Damour, Thibault and Jaranowski, Piotr and Schäfer, Gerhard (2014); Antonelli et al. (2020); Mandal et al. (2023); Kim et al. (2023); Levi and Steinhoff (2015) by finding a suitable canonical transformation to its PN-expanded counterpart. [^3] We thank Mohammed Khalil for providing us with suitable PN Hamiltonians to compare with. The non-spinning component is determined only up to quadratic order in eccentricity ($p_{r}^{2}$) in the tail integral, as higher powers in eccentricity appear at lower PM orders. Thus, we ensured that the 1PM–3PM (tail-free) non-spinning dynamics are unmodified by the presence of the 4PN correction (5). [^4] As a spin-dependent bound-orbit correction would only arise at 5.5PN order, we choose not to include it.
Figure 2: Non-spinning binding energy as a function of the (quasi-circular) velocity $v=(M\dot{\phi})^{2/3}$, for the (calibrated) SEOBNRv5 with $a_{6}$ and (uncalibrated) SEOB-PM Hamiltonians (both along a circular orbit Antonelli et al. (2019); Khalil et al. (2022) and inspiral) across different mass ratios $q=m_{1}/m_{2}$. The shaded region is an estimate of the NR uncertainty Ossokine et al. (2018). The lower panel shows the fractional difference.
Figure 3: Spin-orbit (left-hand panel) and spin-squared (right-hand panel) contributions to the binding energy, for an equal-mass BBH, as a function of the (quasi-circular) velocity $v=(M\dot{\phi})^{2/3}$, for the (calibrated) SEOBNRv5 with ($a_{6}$,$d_{\rm SO}$) and (uncalibrated) SEOB-PM Hamiltonians at different PM orders. The vertical line represents the merger of the NR configuration (the one at the lowest velocity among those used), with the number of GW cycles (top axis) referring to the same simulation. The shaded regions are estimates of the NR uncertainty Ossokine et al. (2018). The lower panel shows the absolute value of the fractional difference. The feature in the lower-right panel around $v\sim 0.4$ is due to a zero-crossing.
Finally, let us comment on the appearance of special functions in $H_{\rm eff}$. Starting at 3PM order we encounter the combination ${{\rm arccosh}(\gamma)}/{\sqrt{\gamma^{2}-1}}$. As ${\rm arccosh}(\gamma)$ and $\sqrt{1-\gamma^{2}}$ are both imaginary valued when $\gamma<1$, we find it convenient to replace this combination by ${\arccos(\gamma)}/{\sqrt{1-\gamma^{2}}}$, which has the same small-velocity expansion for scattering kinematics. At 4PM order we then encounter logarithms, dilogarithms (${\rm Li}_{2}$) and elliptic functions (K/E) of the first and second kind, all evaluated as functions of $\gamma$. In this case, we also find it convenient to introduce the inverse tangent integral ${\rm Ti}_{2}(x):=\int_{0}^{x}({\rm d}t/t)\arctan t$, analogously to what is done above. Fast numerical routines exist for evaluating all of these functions in Cython Behnel et al. (2011), and this leads to an efficient numerical evaluation within pySEOBNR Mihaylov et al. (2023).
Comparing SEOB-PM and NR binding energies during the inspiral – In EOB models one has access to the binary’s dynamics, which enables testing their accuracy by comparing (gauge-invariant) dynamical quantities such as the binding energy Damour et al. (2012); Le Tiec et al. (2012); Nagar et al. (2016); Ossokine et al. (2018); Khalil et al. (2022) and periastron advance Le Tiec et al. (2011); Hinderer et al. (2013). As SEOBNR-PM’s essential new feature is its PM-informed SEOB-PM Hamiltonian, the binding energy is a particularly relevant quantity to compare with NR data. Previous comparisons in the non-spinning case Antonelli et al. (2019); Khalil et al. (2022) have focused on the binding energy computed for circular orbits (i.e., ignoring radiation-reaction effects), although Ref. Antonelli et al. (2019) investigated the effect of neglecting dissipation (see Fig. 6 therein). We instead compute the (dimensionless) binding energy by evaluating ${\cal E}=(H_{\mathrm{EOB}}-M)/\mu$ along the inspiraling dynamics, and compare with NR–binding-energy data from Ref. (Ossokine et al., 2018). Fig. 2 shows the EOB and NR non-spinning binding energies as a function of the (quasi-circular) velocity parameter $v=(M\dot{\phi})^{1/3}$, for SEOBNRv5 with $a_{6}$ and SEOB-PM for circular orbits and along an inspiral. We stress that for the former, a 5PN-unknown parameter ($a_{6}$) in the $A$-potential has been calibrated against 18 non-spinning simulations (see below). Both models show excellent agreement with NR during most of the inspiral, with errors within the NR uncertainty (represented by the gray region) until around 1 GW cycle before merger. The (uncalibrated) SEOB-PM maintains agreement within NR error up to slightly higher velocities for higher mass ratios, and it has much better agreement than when computed on circular orbits Antonelli et al. (2019); Khalil et al. (2022).
We also extract different spin contributions to the binding energy by combining results from NR simulations for various equal-mass spin combinations Dietrich et al. (2017); Ossokine et al. (2018): ${\cal{E}}_{\mathrm{SO}}=-\frac{1}{6}{\cal{E}}(-0.6,0)+\frac{8}{3}{\cal{E}}(0.3,0)-2{\cal{E}}(0,0)-\frac{1}{2}{\cal{E}}(0.6,0)+\mathcal{O}(S^{3})$ and ${\cal{E}}_{\mathrm{S^{2}}}=\frac{3}{2}{\cal{E}}(-0.6,0)-2{\cal{E}}(0,0)+\frac{3}{2}{\cal{E}}(0.6,0)-{\cal{E}}(0.6,-0.6)+\mathcal{O}(S^{3})$, where ${\cal{E}}(\chi_{1},\chi_{2})$ denotes the binding energy in a simulation with dimensionless spins $\chi_{i}$. In Fig. 3 we illustrate the spin-orbit and spin-squared contributions for an equal-mass BBH to the binding energy for the (uncalibrated) SEOB-PM at different PM orders, as compared with NR and with the (calibrated) SEOBNRv5 with ($a_{6}$, $d_{\rm SO}$). Despite not being calibrated to NR, SEOB-PM shows excellent agreement with the NR results, with a clear convergence toward the NR prediction, as more PM corrections are included. Its accuracy is somewhat better than SEOBNRv5, despite the latter model using a Hamiltonian calibrated in the non-spinning ($a_{6}$) and spin-orbit coupling ($d_{\rm SO}$) sector (see below).
Calibration to numerical-relativity waveforms – As discussed, the accuracy of EOB inspiral-merger-ringdown waveforms can be enhanced through calibration to NR simulations. For the inspiral-plunge stage, this is generally achieved by introducing in the Hamiltonian high-order (still unknown) PN terms, whose coefficients are tuned to NR, and fitting the time of merger (i.e., the (2,2)-mode’s peak time) to NR. In the SEOBNRv5 model Pompili et al. (2023), which was built integrating PN results in the Hamiltonian, three calibration parameters were employed: $(\Delta t_{\rm NR},a_{6},d_{\rm SO})$. The parameter $\Delta t_{\rm NR}$ is defined by $t_{\text{peak }}^{22}=t_{\mathrm{ISCO}}+\Delta t_{\mathrm{NR}}$ (see also Eq. (2)), where $t_{\rm{ISCO}}$ is the time at which $r=r_{\rm{ISCO}}$, with $r_{\rm{ISCO}}$ the radius of the Kerr innermost stable circular orbit (ISCO) Bardeen et al. (1972) with the mass and spin of the remnant BH, as given by NR fitting formula  Jiménez-Forteza et al. (2017); Hofmann et al. (2016). The parameter $a_{6}$ is a 5PN correction to the $A$-potential and $d_{\rm SO}$ is a 4.5PN correction in the gyro-gravitomagnetic coefficients. [^5] More specifically, SEOBNRv5’s Hamiltonian employs all known 5PN coefficients in the $A$- and $D$-potentials (except for the partially known 5PN local part of the $A$-potential that is fully replaced by the $a_{6}$ calibration coefficient), and the 5.5PN terms in the $Q$-potential. Although the 4.5PN spin-orbit couplings are known, Ref. Pompili et al. (2023) did not use them (except for the 3.5PN) because the authors showed that having a spin-orbit calibration parameter at 5.5PN (instead of 4.5PN) is not very effective. Here, for the SEOBNR-PM model, we do not calibrate high-order PN terms in the non-spinning and spin sectors of the Hamiltonian (3), but we calibrate only the merger’s time through $\Delta t_{\rm NR}$. In future work, we will explore NR calibrations tailored to the particular structure of the PM terms. Henceforth, we compare the PM-informed model with several versions of the most recent PN-GSF–informed SEOBNRv5, with and without calibration.
Waveform accuracy is often quantified in terms of the mismatch $\mathcal{M}$, defined as $1$ minus the overlap between the normalized waveforms, maximized over a relative time and phase shift:
$\mathcal{M}=1-\max_{\phi_{0},t_{0}}\frac{\left(h_{1}\mid h_{2}\right)}{\sqrt{\left(h_{1}\mid h_{1}\right)\left(h_{2}\mid h_{2}\right)}}.$ (6)
[…]
To calibrate the SEOBNR-PM model, we closely follow the procedure outlined in Refs. Bohé et al. (2017); Pompili et al. (2023); Mihaylov et al. (2023). This procedure essentially consists of determining values of the calibration parameters that minimize a combination of the mismatch and the difference in merger time (defined as the peak of the $(2,2)$-mode amplitude) between EOB and NR waveforms with the same physical parameters ($q,\chi_{1},\chi_{2}$). This is carried out in a Bayesian fashion using the Bilby Ashton et al. (2019) package, and the pySEOBNR code Mihaylov et al. (2023). Finally, we interpolate the best-fit values for each NR simulation across the ($q,\chi_{1},\chi_{2}$) parameter space. As said, in our SEOBNR-PM model, we only calibrate the $\Delta t_{\rm NR}$ parameter (see the Supplemental Material for its expression) using 441 NR simulations of aligned-spin BBHs produced with the pseudo-Spectral Einstein code of the Simulating eXtreme Spacetimes (SXS) Collaboration Boyle et al. (2019); Mroue et al. (2013); Hemberger et al. (2013); Scheel et al. (2015); Lovelace et al. (2015); Chu et al. (2016); Blackman et al. (2015); Abbott et al. (2016c, d); Bohé et al. (2017); Lovelace et al. (2016); Blackman et al. (2017b); Varma et al. (2019b, a); Yoo et al. (2022), which were also employed in Ref. Pompili et al. (2023) for the SEOBNRv5 model. They cover mass ratios $q=m_{1}/m_{2}$ from $1$ to $20$ in the non-spinning limit, and dimensionless spin values going from $-0.998\leq\chi_{i}\leq 0.998$ for $q=1$ to $-0.5\leq\chi_{1}\leq 0.5,~{}\chi_{2}=0$ for $q=15$.
Figure 4: Cumulative maximum mismatch over the binary’s total-mass range $10M_{\odot}\leq M\leq 300M_{\odot}$ for the (calibrated) SEOBNR-PM and SEOBNRv5 models. The study uses 441 SXS NR waveforms, and focuses on the $(\ell,m)=(2,2)$ mode. The vertical dashed lines indicate the medians of the mismatch distributions.
SEOBNR-PM waveform-model performance – To assess the accuracy of the waveform model, we compute its mismatch against the set of 441 SXS NR simulations, and compare its performance to the SEOBNRv5 $(\Delta t_{\rm NR},a_{6},d_{\rm SO})$ model, as well as, to a version of SEOBNRv5 calibrated only via $\Delta t_{\mathrm{NR}}$. Fig. 4 illustrates the cumulative maximum mismatch against the NR simulations over the binary’s total-mass range $10M_{\odot}\leq M\leq 300M_{\odot}$, for the $(\ell,m)=(2,2)$ mode. The overall mismatch of SEOBNR-PM against NR falls roughly between that of the two SEOBNRv5 variations, with a median value $\mathcal{M}_{\mathrm{median}}\sim 6.1\times 10^{-4}$. This represents a remarkably good agreement. When tuning only $\Delta t_{\rm NR}$, we observe that the accuracy of both SEOBNR-PM and SEOBNRv5 tends to degrade for configurations with large positive spins. This results in a tail of high-mismatch cases above $\sim 1\%$, more pronounced for SEOBNRv5, which includes spin-orbit (3.5PN), spin-square (4PN), and spin-cube (3.5PN) effects at a lower PN order than SEOBNR-PM, which includes spin terms up to 5PM order  [^6] When using the SEOBNRv5 model with 4.5PN spin-orbit couplings and calibrate $\Delta t_{\rm NR}$, we find a better agreement than the model with 3.5PN, but SEOBNR-PM performs still better.. Resumming the PM-EOB potentials and introducing calibration parameters could greatly improve SEOBNR-PM’s accuracy for these cases, similar to the calibrated SEOBNRv5. We leave this important work to the future.
Conclusions – In this Letter, we took advantage of the flexible and efficient pySEOBNR code Mihaylov et al. (2023) and recent prediction for the scattering angle in the EOB formalism Buonanno et al. (2024) to build the first inspiral-merger-ringdown EOB waveform model (SEOBNR-PM) for aligned-spin BHs that uses a PM-informed Hamiltonian (i.e., expanded in $G$, but at all orders in the velocity). Importantly, we found that the SEOB-PM non-spinning binding energy, computed along an inspiraling trajectory, at 4PM, and its spin-orbit and spin-spin contributions through 5PM, agree remarkably well with the NR data up to 1 GW cycle before merger (see Figs. 2 and  3). The agreement is comparable and in some cases better than SEOBNRv5, which however was calibrated to NR results Pompili et al. (2023). Furthermore, we calibrated SEOBNR-PM to 441 NR simulations provided by the SXS Collaboration Boyle et al. (2019); Mroue et al. (2013); Hemberger et al. (2013); Scheel et al. (2015); Lovelace et al. (2015); Chu et al. (2016); Blackman et al. (2015); Abbott et al. (2016c, d); Bohé et al. (2017); Lovelace et al. (2016); Blackman et al. (2017b); Varma et al. (2019b, a); Yoo et al. (2022) by tuning the (2,2)-mode’s peak time (i.e., $\Delta t_{\rm NR}$), and found a median mismatch lower than SEOBNRv5, when the latter is similarly calibrated to NR (see Fig. 4). For now, without optimization, the SEOBNR-PM’s evaluation time is an order of magnitude slower than SEOBNRv5.
Considering the recent attention to the two-body gravitational-scattering problem in quantum-field theory, with a slew of new results produced Bern et al. (2021a, 2022a); Damgaard et al. (2023); Dlapa et al. (2022, 2023a); Jakobsen et al. (2023a, 2024), we see the development of the SEOBNR-PM model as a watershed moment — the first true application of these methods to an astrophysically relevant inspiral-merger-ringdown waveform model. Yet, this is only a first step. Given the relevant progress at 5PM Driesse et al. (2024), we hope to incorporate the complete 5PM scattering angle into our effective Hamiltonian in the near future. Recent results separating the local from non-local parts of the 4PM angle Dlapa et al. (2024) will likely be crucial for achieving good agreement with NR for highly elliptic bound systems — ultimately, this may be the SEOBNR-PM model’s raison d’être. In light of the progress in PM fluxes Herrmann et al. (2021a, b); Di Vecchia et al. (2021); Heissenberg (2021); Bjerrum-Bohr et al. (2021); Damgaard et al. (2021); Mougiakakos et al. (2021); Manohar et al. (2022); Dlapa et al. (2023b); Jakobsen et al. (2021, 2022c, 2023b), PM corrections could also be fed into the EOB radiation-reaction forces and gravitational modes. The SEOB-PM Hamiltonian and fluxes will also need to be extended to the astrophysically relevant precessing-spin case. We leave these tantalizing prospects for future work.
Acknowledgments – We are grateful to Zvi Bern, Gustav Uhre Jakobsen, Mohammed Khalil, Jan Plefka and Jan Steinhoff for valuable discussions and comments on this Letter, and to Raffi Enficiaud for his assistance with scientific computing. The work of G.M. and R.P. was supported by the Deutsche Forschungsgemeinschaft (DFG) Projektnummer 417533893/GRK2575 “Rethinking Quantum Field Theory”.
Supplemental Material
Appendix A PM-informed effective Hamiltonian in the EOB approach
Here we present the effective Hamiltonian $H_{\rm eff}$ (3), which depends on the $A$-potential and gyro-gravitomagnetic factors:
$A$ $=\frac{1-2u+\chi_{+}^{2}u^{2}+\Delta A}{1+\chi_{+}^{2}u^{2}(2u+1)}\,,$ $g_{a_{\pm}}$ $=\frac{\Delta g_{a_{\pm}}}{u^{2}}\,.$ (7)
These are PM-expanded in Eq. (4). Corrections to the $A$-potential incorporate even-in-spin PM corrections:
$\Delta A^{(n)}=\sum_{s=0}^{\lfloor(n-1)/2\rfloor}\sum_{i=0}^{2s}\alpha^{(n)}_{(2s-i,i)}\delta^{\sigma(i)}\chi_{+}^{2s-i}\chi_{-}^{i}\,,$ (8)
where $\sigma(i)=0,1$ if $i$ is even or odd, respectively. The gyro-gravitomagnetic factors incorporate odd-in-spin PM corrections:
$\Delta g_{a_{+}}^{(n)}$ $=\!\!\sum_{s=0}^{\lfloor(n-2)/2\rfloor}\sum_{i=0}^{s}\alpha^{(n)}_{(2(s-i)+1,2i)}\chi_{+}^{2(s-i)}\chi_{-}^{2i}\,,$ (9a)
$\Delta g_{a_{-}}^{(n)}$ $=\!\!\sum_{s=0}^{\lfloor(n-2)/2\rfloor}\sum_{i=0}^{s}\alpha^{(n)}_{(2(s-i),2i+1)}\chi_{+}^{2(s-i)}\chi_{-}^{2i}\,.$ (9b)
The dimensionless parameters $\alpha^{(n)}_{(i,j)}$ are functions of $\gamma=E_{\rm eff}/\mu$ and the symmetric mass ratio $\nu$.
We now provide the deformation coefficients required in order to fully specify the effective Hamiltonian for non-spinning configurations. For the full spinning, we refer the interested reader to the ancillary file attached to the arXiv submission of this Letter that contains the complete EOB and effective Hamiltonians. Firstly, the 2PM non-spinning deformation is
$\alpha^{(2)}_{(0,0)}=\frac{3(\Gamma-1)(5\gamma^{2}-1)}{2\gamma^{2}\Gamma}\,,$ (10)
where $\Gamma=\sqrt{1+2\nu(\gamma-1)}$ is the dimensionless total energy. Next, at 3PM order we require
$\alpha^{(3)}_{(0,0)}=\frac{9(\Gamma\!-\!1)\left(30\gamma^{4}\!-\!31\gamma^{2}\!+\!5\right)+2\nu\left(-214\gamma^{5}\!+\!270\gamma^{4}\!+\!323\gamma^{3}\!-\!279\gamma^{2}\!-\!145\gamma\!+\!45\right)}{6\gamma^{2}\left(\gamma^{2}-1\right)\Gamma^{2}}+\frac{4\nu\left(4\gamma^{4}\!-\!12\gamma^{2}\!-\!3\right)}{\gamma^{2}\Gamma^{2}}\frac{\arccos\gamma}{\sqrt{1-\gamma^{2}}}\,.$ (11)
Here, we have replaced ${\rm arccosh}(\gamma)/\sqrt{\gamma^{2}-1}$ by $\arccos(\gamma)/\sqrt{1-\gamma^{2}}$. Finally, at 4PM order we encounter
$\alpha^{(4)}_{(0,0)}=\frac{7\nu\left(380\gamma^{2}+169\right)}{8(\gamma-1)\gamma^{2}\Gamma^{3}}{\rm E}^{2}\left(\frac{\gamma-1}{\gamma+1}\right)+\frac{\left(1200\gamma^{2}+2095\gamma+834\right)\nu}{4\gamma^{2}\left(\gamma^{2}-1\right)\Gamma^{3}}{\rm K}^{2}\left(\frac{\gamma-1}{\gamma+1}\right)$ (12)
$+\frac{\left(-1200\gamma^{3}-2660\gamma^{2}-2929\gamma-1183\right)\nu}{4\gamma^{2}\left(\gamma^{2}-1\right)\Gamma^{3}}{\rm E}\left(\frac{\gamma-1}{\gamma+1}\right){\rm K}\left(\frac{\gamma-1}{\gamma+1}\right)$
$+\frac{\left(-25\gamma^{6}+30\gamma^{4}+111\gamma^{2}+20\right)\nu}{\gamma^{2}\Gamma^{3}}\text{Li}_{2}\left(\frac{1-\gamma}{1+\gamma}\right)+\frac{(\gamma+1)\left(25\gamma^{5}-25\gamma^{4}-5\gamma^{3}+65\gamma^{2}+64\gamma+12\right)\nu}{2\gamma^{2}\Gamma^{3}}\text{Li}_{2}\left(\frac{\gamma-1}{\gamma+1}\right)$
$+\frac{\left(35\gamma^{4}+120\gamma^{3}+90\gamma^{2}+152\gamma+27\right)\nu}{2\gamma^{2}\Gamma^{3}}\log^{2}\left(\frac{\gamma+1}{2}\right)-\frac{4\left(2\gamma^{2}-3\right)\left(15\gamma^{2}-15\gamma+4\right)\nu}{\gamma(\gamma+1)\Gamma^{3}}\frac{{\rm Ti}_{2}\left(\sqrt{\frac{1-\gamma}{1+\gamma}}\right)}{\sqrt{1-\gamma^{2}}}$
$+\frac{\left(2\gamma^{2}-3\right)^{2}\left(35\gamma^{4}-30\gamma^{2}+11\right)\nu}{8\left(\gamma^{2}-1\right)^{3}\Gamma^{3}}\arccos^{2}\gamma+\frac{2\left(75\gamma^{6}-140\gamma^{4}-283\gamma^{2}-852\right)\nu}{3\gamma\left(\gamma^{2}-1\right)\Gamma^{3}}\log(\gamma)$
$+\frac{\left(210\gamma^{6}-552\gamma^{5}+339\gamma^{4}-912\gamma^{3}+3148\gamma^{2}-3336\gamma+1151\right)\nu}{12\gamma^{2}\left(\gamma^{2}-1\right)\Gamma^{3}}\log\left(\frac{u}{4}\right)$
$+\bigg{(}\frac{\left(-35\gamma^{4}-60\gamma^{3}+150\gamma^{2}-76\gamma+5\right)\nu}{2\gamma^{2}\Gamma^{3}}\log\left(\frac{u}{4}\right)$
$\qquad+\frac{\left(-75\gamma^{7}+416\gamma^{5}+612\gamma^{4}+739\gamma^{3}+136\gamma^{2}+2520\gamma+152\right)\nu}{3\gamma^{2}\left(\gamma^{2}-1\right)\Gamma^{3}}\bigg{)}\log\left(\frac{\gamma+1}{2}\right)$
$+\bigg{(}\frac{\left(-420\gamma^{9}+96\gamma^{8}-48\gamma^{7}+5328\gamma^{6}-5279\gamma^{5}-1584\gamma^{4}+7142\gamma^{3}-9360\gamma^{2}+3453\gamma+720\right)\nu}{12\gamma^{2}(\gamma^{2}-1)^{2}\Gamma^{3}}$
$\qquad-\frac{48\left(7\gamma^{2}-5\right)\left(4\gamma^{4}-12\gamma^{2}-3\right)(\Gamma-1)\nu}{12\gamma^{2}(\gamma^{2}-1)\Gamma^{3}}$
$\qquad-\frac{\left(2\gamma^{2}-3\right)\left(35\gamma^{4}-30\gamma^{2}+11\right)\nu}{4\gamma(1-\gamma^{2})\Gamma^{3}}\log\left(\frac{u}{4}\right)+\frac{4\left(2\gamma^{2}-3\right)\left(15\gamma^{2}+2\right)\nu}{\gamma(1-\gamma^{2})\Gamma^{3}}\log\left(\frac{\gamma+1}{2}\right)\bigg{)}\frac{\arccos\gamma}{\sqrt{1-\gamma^{2}}}$
$+(\Gamma-1)\bigg{(}\frac{5115\gamma^{8}-9537\gamma^{6}+5657\gamma^{4}-1115\gamma^{2}+72}{16\gamma^{4}\left(\gamma^{2}-1\right)^{2}\Gamma^{3}}$
$\qquad+\frac{\left(8159\gamma^{8}-3136\gamma^{7}-23601\gamma^{6}-3360\gamma^{5}+15409\gamma^{4}+4000\gamma^{3}-1995\gamma^{2}+108\right)\nu}{24(\gamma-1)\gamma^{4}(\gamma+1)^{2}\Gamma^{3}}\bigg{)}$
$+\frac{\nu}{144\gamma^{9}\left(\gamma^{2}-1\right)^{2}\Gamma^{3}}\bigg{(}-600\pi^{2}\gamma^{17}+3600\gamma^{16}+480\left(9+4\pi^{2}\right)\gamma^{15}+2\left(720\pi^{2}-28843\right)\gamma^{14}+\left(36759-5136\pi^{2}\right)\gamma^{13}$
$\qquad+\left(44698-1056\pi^{2}\right)\gamma^{12}+\left(6624\pi^{2}-43235\right)\gamma^{11}+\left(7702-2208\pi^{2}\right)\gamma^{10}-5\left(2155+504\pi^{2}\right)\gamma^{9}$
$\qquad+2\left(23947+912\pi^{2}\right)\gamma^{8}-\left(45605+288\pi^{2}\right)\gamma^{7}+12701\gamma^{6}+648\gamma^{5}-1471\gamma^{4}+207\gamma^{2}-45\bigg{)}\,,$
where ${\rm Ti}_{2}(x):=\int_{0}^{x}({\rm d}t/t)\arctan t$ is the inverse tangent integral. This expression also includes $\log(u)$, which is the logarithmic bound-orbit counterpart to $\log(\gamma^{2}-1)$. The $\log(u)$ can be identified by taking the bound circular-orbit limit (at leading-PN order) of $\log(1-\gamma^{2})$, i.e. the real part of $\log(\gamma^{2}-1)$, ensuring that the coefficient is real for $\gamma<1$. Finally, to complete the non-spinning dynamics we require the 4PN correction coefficient $\Delta A^{\rm 4PN}$ (5). This includes three numerical coefficients:
$c_{1}=\frac{\nu}{15}(-1411\!+\!296\gamma_{\rm E}\!-\!1328\log 2\!+\!2187\log 3)\,,$ (13a)
$c_{2}=\frac{9\nu^{3}}{4}\!+\!\frac{\nu^{2}}{192}\left(615\pi^{2}-16408\right)$ (13b)
$\!+\!\nu\left(-\frac{51187}{360}\!+\!\frac{136\gamma_{\rm E}}{3}\!+\!\frac{1571\pi^{2}}{6144}-\frac{856\log 2}{15}\!+\!\frac{729\log 3}{5}\right)\!,$
$c_{3}=-\frac{68\nu}{3}\,,$ (13c)
with $\gamma_{\rm E}\approx 0.557$ Euler’s constant.
Appendix B Interpreting $\gamma$ within $\alpha^{(n)}_{(i,j)}$
The effective Hamiltonian $H_{\rm eff}$ (3) contains deformation parameters $\alpha^{(n)}_{(i,j)}$ that depend on $\gamma=E_{\rm eff}/\mu$ (i.e., the Hamiltonian technically depends on itself). In principle, one should therefore solve for $E_{\rm eff}=H_{\rm eff}$ as a function of the kinematic variables ($r$,$p_{r}$,$p_{\phi}$); however, given the highly non-trivial dependence on $\gamma$ on the right-hand side of Eq. (3), not to mention the appearance of special functions including dilogarithms and elliptics, such an approach is not practical. While one might also try solving for the Hamiltonian numerically at a given phase-space point, this approach will not lead to an efficient implementation within pySEOBNR.
Instead, building on the approach taken in Ref. Antonelli et al. (2019); Khalil et al. (2022) for the non-spinning case, we interpret $\gamma$ as the effective energy expanded only up to whatever PM and/or spin order is necessary to ensure that the complete resummed Hamiltonian $H_{\rm EOB}=M\,\sqrt{1+2\nu(H_{\rm eff}/\mu-1)}$ is consistent with the known PM and spin results. The effective Hamiltonian may be perturbatively expanded as
$\gamma=\gamma_{\rm Kerr}+\sum_{n\geq 2}\sum_{s\geq 0}\Delta_{(s)}^{(n)}(\gamma_{\rm Kerr})$ (14)
with deformations $\Delta_{(s)}^{(n)}(\gamma_{\rm Kerr})$, characterized by PM order $n$ and spin order $s$, depending on $\gamma_{\rm Kerr}=H_{\rm Kerr}/\mu$. The Kerr Hamiltonian here is
$H_{\rm Kerr}=\frac{2Mp_{\phi}a}{r^{3}+a^{2}(r+2M)}$ (15)
$\quad+\sqrt{A^{\rm Kerr}\left(\mu^{2}+\frac{p_{\phi}^{2}}{r^{2}}+(1+B_{\rm np}^{\rm Kerr})p_{r}^{2}+B_{\rm npa}^{\rm Kerr}\frac{p_{\phi}^{2}a^{2}}{r^{2}}\right)}\,,$
where we identify the total spin $a=a_{+}$ in the EOB. The three functions appearing in the Kerr Hamiltonian are
$A^{\rm Kerr}$ $=\frac{1-2u+\chi^{2}u^{2}}{1+\chi^{2}u^{2}(2u+1)}\,,$ (16a)
$B_{\rm np}^{\rm Kerr}$ $=\chi^{2}u^{2}-2u\,,$ (16b)
$B_{\rm npa}^{\rm Kerr}$ $=-\frac{1+2u}{r^{2}+a^{2}(1+2u)}\,.$ (16c)
For our purposes, expressions for the following five deformations (up to 3PM order) are sufficient:
$\Delta_{(0)}^{(2)}$ $=\frac{u^{2}}{2}\alpha^{(2)}_{(0,0)}(\gamma_{\rm Kerr})\gamma_{\rm Kerr}\,,$ (17a)
$\Delta_{(1)}^{(2)}$ $=\ell u^{3}\left((\alpha^{(2)}_{(1,0)}(\gamma_{\rm Kerr})-2)a_{+}+\alpha^{(2)}_{(0,1)}(\gamma_{\rm Kerr})\delta a_{-}\right),$ (17b)
$\Delta_{(0)}^{(3)}$ $=\frac{u^{3}}{2}\left(2\alpha^{(2)}_{(0,0)}(\gamma_{\rm Kerr})+\alpha^{(3)}_{(0,0)}(\gamma_{\rm Kerr})\right)\gamma_{\rm Kerr}\,,$ (17c)
$\Delta_{(1)}^{(3)}$ $=\ell u^{4}\left(\alpha^{(3)}_{(1,0)}(\gamma_{\rm Kerr})a_{+}+\alpha^{(3)}_{(0,1)}(\gamma_{\rm Kerr})\delta a_{-}\right)\!,$ (17d)
$\Delta_{(2)}^{(3)}$ $=\frac{u^{3}}{2}\left(\alpha^{(3)}_{(2,0)}(\gamma_{\rm Kerr})a_{+}^{2}+\alpha^{(3)}_{(0,2)}(\gamma_{\rm Kerr})a_{-}^{2}\right.$
$\qquad\qquad\qquad+\left.\alpha^{(3)}_{(1,1)}(\gamma_{\rm Kerr})\delta a_{+}a_{-}\right)\gamma_{\rm Kerr}\,,$ (17e)
Within a given coefficient $\alpha_{(i,j)}^{(n)}$ we use whatever deformations in Eq. (17) are required. The explicit coefficient-by-coefficient replacements used in our model are
$\alpha^{(2)}_{(0,0)}(\gamma)$ ::: $\gamma$ $\to\gamma_{\rm Kerr}+\Delta^{(2)}_{(0)}+\Delta^{(2)}_{(1)}+\Delta^{(3)}_{(1)}+\Delta^{(3)}_{(2)}\,,$
$\alpha^{(3)}_{(0,0)}(\gamma)$ ::: $\gamma$ $\to\gamma_{\rm Kerr}+\Delta^{(2)}_{(1)}\,,$
$\alpha^{(2)}_{(1,0)}(\gamma)$ ::: $\gamma$ $\to\gamma_{\rm Kerr}+\Delta^{(2)}_{(0)}\!+\!\Delta^{(2)}_{(1)}\!+\!\Delta^{(3)}_{(0)}\!+\!\Delta^{(3)}_{(1)}\!+\!\Delta^{(3)}_{(2)}\,,$
$\alpha^{(2)}_{(0,1)}(\gamma)$ ::: $\gamma$ $\to\gamma_{\rm Kerr}+\Delta^{(2)}_{(0)}\!+\!\Delta^{(2)}_{(1)}\!+\!\Delta^{(3)}_{(0)}\!+\!\Delta^{(3)}_{(1)}\!+\!\Delta^{(3)}_{(2)}\,,$
$\alpha^{(3)}_{(1,0)}(\gamma)$ ::: $\gamma$ $\to\gamma_{\rm Kerr}+\Delta^{(2)}_{(0)}+\Delta^{(2)}_{(1)}\,,$
$\alpha^{(3)}_{(0,1)}(\gamma)$ ::: $\gamma$ $\to\gamma_{\rm Kerr}+\Delta^{(2)}_{(0)}+\Delta^{(2)}_{(1)}\,,$
$\alpha^{(3)}_{(2,0)}(\gamma)$ ::: $\gamma$ $\to\gamma_{\rm Kerr}+\Delta^{(2)}_{(0)}+\Delta^{(2)}_{(1)}\,,$
$\alpha^{(3)}_{(0,2)}(\gamma)$ ::: $\gamma$ $\to\gamma_{\rm Kerr}+\Delta^{(2)}_{(0)}+\Delta^{(2)}_{(1)}\,,$
$\alpha^{(3)}_{(1,1)}(\gamma)$ ::: $\gamma$ $\to\gamma_{\rm Kerr}+\Delta^{(2)}_{(0)}+\Delta^{(2)}_{(1)}\,,$ (18)
and in all other cases we replace $\gamma\to\gamma_{\rm Kerr}$ (including $\Delta A^{\rm 4PN}$). For example, as $\alpha^{(3)}_{(1,0)}$ appears at 3PM order in $H_{\rm eff}$, we need all corrections up to 2PM order in order to ensure that our Hamiltonian remains correct up to 5PM. An exception is the two non-spinning parameters $\alpha^{(2)}_{(0,0)}$ and $\alpha^{(3)}_{(0,0)}$, wherein we do not include $\Delta^{(3)}_{(0)}$ and $\Delta^{(2)}_{(0)}$, respectively, as the non-spinning 5PM contribution is not included from known perturbative PM results. We instead correct using the 4PN term $\Delta A^{\rm 4PN}$. The result of this procedure is an effective Hamiltonian that now depends explicitly on ($r$,$p_{r}$,$p_{\phi}$). As this procedure changes the nature of the resummation, our Hamiltonian is therefore different from the one encountered in the SEOB-PM scattering model Buonanno et al. (2024).
A consequence of treating $\gamma$ via Eq. (14) is that it vanishes wherever $H_{\rm Kerr}$ does. For example, in the non-spinning limit, this leads to a pole at the Schwarzschild horizon $r=2M$ in the $\alpha^{(n)}_{(i,j)}$ coefficients (8), which contain powers of $1/\gamma$, and consequently in the $A$-potential. This would not necessarily occur if $\gamma$ was defined as $\gamma=E_{\rm eff}/\mu$, as the effective Hamiltonian does not need to vanish at the horizon. Since the pole appears at a very small separation, it does not prevent evolving the binary until the merger time (which occurs before the time of horizon crossing) and obtaining an accurate inspiral-merger-ringdown waveform. However, the presence of the pole impacts the shape of the $A$-potential (and of the overall Hamiltonian) in the strong-field regime and might limit the flexibility that can be gained from incorporating higher-order calibration parameters in the EOB Hamiltonian. For example, adding a 5PN calibration parameter $a_{6}u^{6}$ to the $A$-potential via Eq. (5), similar to the one used in SEOBNRv5, does not improve the agreement of the non-spinning SEOBNR-PM against NR simulations as effectively as the corresponding parameter in SEOBNRv5. We recall that SEOBNRv5 is highly faithful to NR simulations in the non-spinning limit, with a median mismatch of $4.6\times 10^{-5}$ and only $2$ cases having a mismatch $\gtrsim 1\times 10^{-4}$, out of 40 SXS NR waveforms. In the future, we will explore different strategies to interpret $\gamma$, and seek for a more suitable choice of the NR calibration parameters tailored to the particular structure of the PM terms, which can enable a more effective NR calibration of the spinning model.
Figure 5: Mismatch over the binary’s total-mass range $10M_{\odot}\leq M\leq 300M_{\odot}$ for the SEOBNR-PM and SEOBNRv5 models calibrated using only the $\Delta t_{\rm{NR}}$ parameter. The study uses 441 SXS NR waveforms, and focuses on the $(\ell,m)=(2,2)$ mode. Spinning and non-spinning configurations are shown in blue and orange respectively.
Appendix C Fits of the NR-calibration parameters
Since NR simulations are only available at discrete parameter values, to generate the model for generic configurations we need to fit the NR calibration parameter $\Delta t_{\mathrm{NR}}$ across the ($q,\chi_{1},\chi_{2}$), or equivalently ($\nu,\chi_{+},\chi_{-}$), parameter space. We do this hierarchically, fitting first non-spinning and then aligned-spin configurations, using the same ansatz as in SEOBNRv5. The final expression for SEOBNR-PM reads
$\Delta t_{\mathrm{NR}}$ $=\nu^{-1/5+12.73\nu}\big{(}113115.96\nu^{3}-25626.22\nu^{2}$ (19)
$\quad-1457.38\nu-60.17\big{)}$
$+\nu^{-1/5}\big{(}195.45\nu^{2}\chi_{+}-190.15\nu^{2}\chi_{-}+52.14\nu\chi_{+}^{2}$
$\quad-72.80\nu\chi_{+}\chi_{-}-72.56\nu\chi_{+}+50.90\nu\chi_{-}^{2}$
$\quad-15.76\nu\chi_{-}+0.24\chi_{+}^{4}-1.51\chi_{+}^{3}+4.23\chi_{+}^{2}\chi_{-}$
$\quad-10.51\chi_{+}^{2}+1.22\chi_{+}\chi_{-}^{2}+18.89\chi_{+}\chi_{-}$
$\quad+10.10\chi_{+}-10.50\chi_{-}^{2}+17.08\chi_{-}\big{)}\,.$
For completeness, we also report the fit used in the variation of the SEOBNRv5 model shown in Fig. 4, in which only this parameter is tuned to NR:
$\Delta t_{\mathrm{NR}}$ $=\nu^{-1/5+4.97\nu}\big{(}-4091.55\nu^{3}+2493.06\nu^{2}$ (20)
$\quad-205.61\nu-53.99\big{)}$
$+\nu^{-1/5}\big{(}45.32\nu^{2}\chi_{+}-874.81\nu^{2}\chi_{-}+71.25\nu\chi_{+}^{2}$
$\quad+65.96\nu\chi_{+}\chi_{-}+3.72\nu\chi_{+}-100.55\nu\chi_{-}^{2}$
$\quad+160.60\nu\chi_{-}-89.85\chi_{+}^{4}-43.49\chi_{+}^{3}-1.07\chi_{+}^{2}\chi_{-}$
$\quad+59.49\chi_{+}^{2}+0.94\chi_{+}\chi_{-}^{2}-17.51\chi_{+}\chi_{-}$
$\quad+36.27\chi_{+}+26.08\chi_{-}^{2}+15.37\chi_{-}\big{)}\,.$
The complete SEOBNRv5 model also contains calibrations of the $a_{6}$ and $d_{\rm SO}$ parameters, together with a different fit for $\Delta t_{\mathrm{NR}}$ Pompili et al. (2023).
Appendix D Mismatch in terms of the total mass
In Fig. 5 we show the mismatch over the binary’s total mass range $10M_{\odot}\leq M\leq 300M_{\odot}$ for the SEOBNR-PM and SEOBNRv5 models calibrated using only the $\Delta t_{\rm{NR}}$ parameter, against the 441 SXS NR waveforms used in this work. Spinning and non-spinning configurations are shown in blue and orange, respectively. This provides a complementary picture to Fig. 4, which only presents the maximum mismatch across the total mass range for each configuration.